The Velocity Distribution for flow over a flat plate is given by u=(2/3)y-y2, Where u is the point velocity in meter per second at a distance y meter above the plate.

Determine the shear stress at y=0 and y=15 cm. Assume dynamic viscosity as 8.63 poises


1
Expert's answer
2021-04-29T07:57:46-0400

Given "u=\dfrac{2}{3}y-y^2." Then


"\dfrac{du}{dy}=\dfrac{2}{3}-2y"


Shear stress is given


"\tau=\mu\cdot\dfrac{du}{dy}"

Assume dynamic viscosity as "\mu=8.63\ \text{poise}=0.863\ N\cdot s/m^2"


At "y=0"

"\dfrac{du}{dy}=\dfrac{2}{3}-2(0)=\dfrac{2}{3}(s^{-1})"


Then


"\tau=0.863\ N\cdot s/m^2\cdot\dfrac{2}{3}s^{-1}=0.575 N/m^2"

The shear stress at y = 0 is "0.575N\/m^2."


At "y=15 cm=0.15m"

"\dfrac{du}{dy}=\dfrac{2}{3}-2(0.15)=\dfrac{1.1}{3}(s^{-1})"


Then


"\tau=0.863\ N\cdot s/m^2\cdot\dfrac{1.1}{3}s^{-1}=0.316 N/m^2"

The shear stress at y = 15 cm is "0.316N\/m^2."